# Transfer learning via Regularized Linear Discriminant Analysis

### v2

@@
Transfer learning via Regularized Linear Discriminant Analysis
Hongzhe Zhang, Arnab Auddy, Hongzhe Lee
(January 7, 2025)
@@ 1 Introduction
The performance of supervised statistical learning methods often relies on the availability of sufficient training data. When data is scarce, choosing an appropriately flexible model becomes critical to achieving optimal prediction accuracy. This is a classic illustration of the well-known “bias-variance tradeoff.” On the other hand, when constructing a predictive model for a specific “target” population, multiple auxiliary “source” datasets may provide valuable supplementary information. Techniques from the field of transfer learning (Pan and Yang, 2009; Weiss et al., 2016; Cai and Wei, 2021; Auddy et al., 2024) aim to exploit these auxiliary datasets to improve the model’s performance for the target population. Other relevant approaches include meta-learning (Peng, 2020; Huisman et al., 2021), domain adaptation (Redko et al., 2020; Sun et al., 2015), and, more recently, continual learning (De Lange et al., 2021).
@@
In this paper, we propose a transfer learning framework for regularized discriminant analysis (RDA) (Friedman, 1989), which we term TL-RDA. Our model assumes a random classification weights setup, where the means of the covariates between two classes differ by a random quantity, $\delta$, with zero mean and constant variance. The auxiliary and target populations are related through the correlation structure of $\delta$. The TL-RDA framework aims to estimate the Bayes optimal predictor by combining naive RDA estimates from both the auxiliary and target populations through a weighted summation. The weights are designed to minimize the distance between the TL-RDA estimator and the Bayes optimal direction in a high-dimensional setting, where the number of features $p$ grows proportionally to the sample size $n$ in all populations. We derive the explicit asymptotic error rate for TL-RDA and show that it achieves the lowest error rate among all estimators based on weighted summations, including naive RDA using only target population data.
In the remainder of this introduction, we provide a detailed problem setup and outline the proposed TL-RDA approach. We introduce various types of weights used to integrate auxiliary information, followed by the technical assumptions required for analyzing the estimator in the context of random matrix theory. Detailed results can be found in subsequent sections.
1.1 Problem Formulation
We begin with formulating the problem.
Assumption 1.1 (Two-class Gaussian).
$(y_{k})_{i}\in\{-1,+1\}\ \ \ \mathbb{P}((y_{k})_{i}=\pm 1)=\pi_{\pm 1}\ \ \ (X_{k})_{i}|(y_{k})_{i}\sim N(\mu_{(y_{k})_{i},k},\Sigma).$ (1)
Here $(X_{k})_{i},i=1,\cdots,n_{k}$ is a $p$-dimensional vector, and for ease of notation, we write $\mathbf{X}_{k}=((X_{k})_{1}\,(X_{k})_{2}\,\ldots(X_{k})_{n_{k}})^{\top}\in\mathbb{R}^{n_{k}\times p}$ as a $n_{k}\times p$ dimensional matrix. We now point out additional details about the parameters of this mixture distribution.
@@ 1.1 Problem Formulation
Note that the mixing proportions $\pi_{\pm 1}$ remain the same across populations. For ease of presentation, we mainly discuss the simpler case $\pi_{-1}=\pi_{+1}=1/2$, and in fact, the general case of unequal proportions can be managed in a similar manner without significant technical novelty (Appendix C). Moreover, in most of the paper, we assume the population covariance matrix are the same for all $K$ populations, and are denoted by $\Sigma$. This assumption is relaxed in Section 4 where we allow each population to have their own covariance matrix $\Sigma_{k}$ for $k=1,\dots,K$. Under this set up, the Bayes optimal prediction direction for the target population $K$ is given by (Anderson, 1958), where
$d_{Bayes}:=\Sigma^{-1}\delta_{K}\quad\text{where}\quad\bar{\mu}_{K}=\frac{\mu_{+1,K}+\mu_{-1,K}}{2}\quad\text{for}\quad\mu_{\pm 1,K}=\bar{\mu}_{K}\pm\delta_{K}.$
@@
The Bayes prediction for a testing data point $x_{0}$ from population $K$ is
@@
where $x_{0}\in\mathbb{R}^{p}$ is a test point. We call this TL-RDA. We use the regular intercept term $\widehat{b}_{K}$ as all intercept terms are asymptotically zero in the regime considered. We formulate two criteria to be optimized over $\mathbf{w}$. Let us recall the Bayes classification direction $d_{Bayes}$. Now for a testing data pair $(x_{0},y_{0})$ sampled from population $K$, we consider the two errors:
@@
In addition, we assume that the population covariance matrix are the same for all $K$ populations (assumption 1.1). With this in mind, a natural way to further exploit mutual information across populations is to use a pooled sample covariance for all discriminant directions $\widehat{d}_{k}$. We write $\widehat{\Sigma}_{P}$ to denote the pooled sample covariance matrix
@@
This form of TL estimator (2) has been used in Zhang and Li (2023) and Dobriban and Sheng (2020) for regression methods. (Helm et al., 2024) also considers a TL estimator that aggregates auxiliary information by weighted summations of individual discriminant directions. However, their aggregation is less adaptive in the form of $\widehat{d}(\mathbf{w})^{*}=a\widehat{d}_{1}+(1-a)\sum_{k=2}^{K}\widehat{d}_{k}/(K-1)$ where $a$ is a constant. In addition, the final weight $a$ minimizes only empirical quantities and provides no guarantee of minimizing the population level criteria such as (3) or (4).
1.2 Technical Details
In this section, we introduce some necessary technical details and pose several assumptions on the population parameters. The analysis of the proposed transfer learning estimator rely heavily on the recent advances in random matrix theory, and its relevant key results are summarized below.
1.2.1 Random Matrix Theory
The sample covariance matrix is an important part of the transfer learning estimator. We consider the Marchenko-Pastur type sample covariance matrices as in several relevant works (Dobriban and Wager, 2018; Tripuraneni et al., 2021; Zhao et al., 2023; Zhang and Li, 2023; Song et al., 2024).
For a symmetric matrix, we can characterize its spectral distribution by a cumulative distribution function that places equal point mass on its eigenvalues. Our asymptotic analysis is based on the convergences of spectral distributions of sample covariance matrices, for which the following assumptions are required.
@@ 1.2.1 Random Matrix Theory
1.2.2 Random Classification Weights
Another key parameter is the classification weights $\delta_{k}$ that separate the two classes. In this paper, we consider a random classification weights set up formalized below
Assumption 1.3 (Random Classification Weights).
The following conditions hold for populations $k=1,\cdots,K$.
  1. 1.
    for fixed $C,\eta>0$.
  2. 2.
    $\bar{\mu}_{k}$ are either fixed or randomly distributed independent of $\delta_{k},\mathbf{Z}_{k},y_{k}$. The second moment of $\bar{\mu}_{k}$ are bounded almost surely such that $\lim\sup_{p\rightarrow\infty}\|\bar{\mu}_{k}\|/p^{1/2-\xi}\leq C$ for some constants $\xi,C>0$.
Assumption 1.4 (Correlated Classification Weights).
The rest of the paper is organized as follows. In Section 2, we present the limiting expressions of the asymptotic prediction error for the proposed TL-RDA / TLP-RDA estimators as well as the optimal weights. Next, in Section 3, we provide guidance on how to choose between population specific transfer learning (TL-RDA) and the pooled covariance based transfer learning (TLP-RDA), as well as weights that minimize either criteria (3) or (4). In Section 4, we relax Assumption 1.2 to allow $\mathbf{X}_{k}-\mathbb{E}(\mathbf{X}_{k})=\mathbf{Z}_{k}\Sigma_{k}^{1/2}$ for unequal covariance matrices $\Sigma_{k}$ for $k=1,\dots,K$. Finally, in Section 5 we implement the proposed transfer learning classifiers on a MESA protein data to predict a binary survival outcome, which demonstrates that the proposed methods outperform naive methods.
2 Asymptotic Expressions
@@ 2.1 Classification error
Under the two-class Gaussian classification model (Assumption 1.1), the expected test error of the linear classifier TL-RDA under weight $\mathbf{w}$ in target population $K$ can be written as
@@
Suppose that assumptions 1.1-1.4 as well as 2.1 and 2.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, we have that the limiting form of $Err(\mathbf{w})$ for a given weight vector $\mathbf{w}\in\mathbb{R}^{K}$ is given by:
@@
The limits of the off-diagonal terms in $\mathcal{A}$ depends on the interplay of sample covariances in separated populations. Similar terms also appear later in the paper when finding the optimal weight vector (see Theorems 2.3, 2.7). Depending on the specific scenario, the term $\mathcal{M}_{kk^{\prime}}$ has explicit expressions, some of which are presented separately in Lemma 5.
@@ 2.2 Minimum Estimation Risk Weight
Suppose that assumptions 1.1-1.4 as well as 2.1 and 2.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, the weight for minimizing the error in estimating the Bayes optimal discriminator $d_{Bayes}$ is given by:
@@ 2.3 Minimum Prediction Risk Weight
Suppose that Assumptions 1.1-1.4 as well as 2.1 and 2.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, the weight for minimizing the excess risk, i.e., the error in predicting the class at a random test point $x_{0}$, when compared to the Bayes optimal discriminator $d_{Bayes}$, is given by:
@@ 2.4 Geometric Interpretation
$d_{est}:=\widehat{d}(\mathbf{w}_{E})\ \ \ d_{err}:=\widehat{d}(\mathbf{w}_{P}).$
Also recall the Bayes discriminant direction $d_{Bayes}$. These discriminant directions are visualized in Figure 1. The blue plane is the space spanned by linear combinations of the local discriminant directions $\{\widehat{d}_{k}\}$, and the top black line stands for the Bayes direction, which is not necessarily in the linear span of $\{\widehat{d}_{k}\}$. Since the TL-RDA directions $d_{est}$ and $d_{err}$ are both linear combinations of $\widehat{d}_{k}$, they both lie on the blue plane and are denoted by colored lines. The direction obtained by minimizing the error in estimating $d_{Bayes}$ is given by $d_{est}$, which is the projection of $d_{Bayes}$ onto the blue plane ${\rm span}\{\widehat{d}_{k}\}$. The criteria 3 suggests $d_{est}$ can be seen as a OLS loss fitting to $d_{Bayes}$ with $d_{k}$.
@@
Since $\Phi^{-1}(\cdot)$ is a monotonically increasing function, it is clear that $\cos(\theta)$ is close to one, i.e., $\theta$ is close to zero, if and only if $Err(\mathbf{w})$ is close to the Bayes error $Err_{Bayes}$. That is, the size of $\theta$ directly quantify the inefficiency of TL-RDA relative to $d_{Bayes}$ in terms of classification error: a smaller $\theta$ is equivalent to a near-optimal transfer combination weight $\mathbf{w}$.
Finally, let us also consider an observation $x_{0}$ pointing to a random direction, and denote the angle between $d_{err}$ and $x_{0}$ as $\beta$ and the angle between $d_{Bayes}$ and $x_{0}$ as $\alpha$. Straightforwardly, $d_{err}$ minimizes the difference between $\beta$ and $\alpha$ as $\mathbf{w}^{P}$ minimizes the difference between the inner products $d_{err}^{\top}x_{0}$ and $d_{Bayes}^{\top}x_{0}$. Proposition 2.11 suggests $d_{err}$ minimizes $\theta$ simultaneously.
@@
3 Weight Selection
In this section, we present some guidance on how one can choose between different weighting schemes. As we have repeatedly illustrated in the previous section, the $d_{est}$ is dominated by $d_{err}$ in terms of classification error on unseen data. This claim, however, holds only when the test data distribution is as given by Assumption 1.1. As we shall demonstrate, $d_{est}$ out performs $d_{err}$ considerably when there is a distribution shift in terms of classification error. The TLP-LDA estimator instead consider the space spanned by $\widehat{d}_{k}^{P}$, thus, the final TL estimator is in a different but related column space. We will prove, at least in special cases, optimal TLP-RDA outperforms the optimal TL-RDA when the aspect ratio is large enough.

### v3

# Transfer Learning via Regularized Random-effects Linear Discriminant Analysis

@@
Abstract
Linear discriminant analysis is a widely used method for classification. However, the high dimensionality of predictors combined with small sample sizes often results in large classification errors. To address this challenge, it is crucial to leverage data from related source models to enhance the classification performance of a target model. This paper proposes a transfer learning approach via regularized random-effects linear discriminant analysis, where the discriminant direction is estimated as a weighted combination of ridge estimates obtained from both the target and source models. Multiple strategies for determining these weights are introduced and evaluated, including one that minimizes the estimation risk of the discriminant vector and another that minimizes the classification error. Utilizing results from random matrix theory, we explicitly derive the asymptotic values of these weights and the associated classification error rates in the high-dimensional setting, where $p/n\rightarrow\gamma$, with $p$ representing the predictor dimension and $n$ the sample size. Extensive numerical studies, including simulations and analysis of proteomics-based 10-year cardiovascular disease risk classification, demonstrate the effectiveness of the proposed approach.
@@ 1 Introduction
Large and diverse data sets are ubiquitous in modern applications, including those in genomics and medical decisions. It is of significant interest to integrate different data sets to obtain more accurate parameter estimates or to make a more accurate prediction or classification of an outcome. The success of a supervised statistical learning method relies on the availability of training data. When data is scarce, choosing an appropriately flexible model becomes critical to achieving optimal prediction accuracy. This is a classic illustration of the well-known “bias-variance tradeoff”. When building a prediction model for a target population, many auxiliary source data sets may exist and provide additional information for building the model for the target population. Modern techniques in the field of transfer learning (Pan and Yang, 2009; Weiss et al., 2016) aim to exploit these additional information. Given a target problem to solve, transfer learning (Torrey and Shavlik, 2010) aims at transferring the knowledge from different but related samples or studies to improve the learning performance of the target problem. In biomedical studies, some clinical or biological outcomes are hard to obtain due to ethical or cost issues, in which case transfer learning can be leveraged to boost the prediction and estimation performance by effectively utilizing information from related studies. Other relevant approaches include meta-learning (Peng, 2020; Huisman et al., 2021), domain adaptation (Redko et al., 2020; Sun et al., 2015), and, more recently, continual learning (De Lange et al., 2021).
In the high dimensional setting, Li et al. (2022, 2023) developed transfer learning methods for sparse high dimensional regressions and demonstrated that one can improve predictions of gene expression levels using data across different tissues. Such sparse models work well when the true models are sparse and the sample sizes are large. However, there are settings where sparse model assumption may not be valid. In genetics, estimating polygenetic risk scores (prss) using genome-wide genotype data (Mak et al., 2017; Torkamani et al., 2018) is an active area of research. Such prss can be used in risk stratification, or can be treated as risk factors in population health studies. However, due to very large number of genetic variants but relatively small sample sizes, building a prs model that accurately predicts the prs scores is challenging. Alternatively, ridge regression, which does not require the sparseness assumption, but can handle the linkage disequilibrium among the genetic variants, provides a viable method for prs prediction. Zhang and Li (2023) studied the estimation and prediction of random coefficient ridge regression in the setting of transfer learning, where in addition to observations from the target model, source samples from different but possibly related regression models are available. The informativeness of the source model to the target model can be quantified by the correlation between the regression coefficients.
@@
The method of Zhang and Li (2023) is developed for continuous outcomes. In this paper, we propose a transfer learning framework for regularized linear discriminant analysis (RDA) (Friedman, 1989), which we term TL-RDA. Our model assumes a random classification weights setup, where the means of the covariates between two classes differ by a random quantity, $\delta$, with zero mean and constant variance. The auxiliary and target populations are related through the correlation structure of $\delta$. The TL-RDA framework aims to estimate the Bayes optimal predictor by combining naive RDA estimates from both the auxiliary and target populations through a weighted summation. The weights are designed to minimize the distance between the TL-RDA estimator and the Bayes optimal direction in a high-dimensional setting, where the number of features $p$ grows proportionally to the sample size $n$ in all populations. We derive the explicit asymptotic error rate for TL-RDA and show that it achieves the lowest error rate among all estimators based on weighted summations, including naive RDA using only target population data.
The remainder of the paper is organized as follows. We first provide a detailed problem setup and outline the proposed TL-RDA approach in Section 2. We then introduce various types of weights used to integrate auxiliary information, followed by the technical assumptions required for analyzing the estimator in the context of random matrix theory in Section 3. In Section 4, we present the analysis of TL-RDA in the proportional regime, providing explicit asymptotic expressions for the different weighting schemes, along with their corresponding error rates. Section 5 offers interpretations of these weights and guidance on how users can select the most appropriate weights for their applications. In the first five sections, we have assumed all populations share the same population covariance matrix. Section 6 extends the TL-RDA to a heterogeneous population covariance matrix set up. Finally, in Section 7, we demonstrate the performance of TL-RDA on a binary 10-year cardiovascular disease risk classification proteomics data in the Chronic Renal Insufficiency Cohort (CRIC).
2 Transfer Learning via Regularized Discriminant Analysis
We consider the setting of two-class LDA in the setting of transfer learning, where we have data observed from both the target model, indexd by $K$, and $K-1$ source models, indexed by $k=1,\cdots,K-1$. We assume that all models, $k=1,\cdots,K$, follow the classic two-class Gaussian mixture model. More specifically, for $i=1,\dots,n_{k}$ and $k=1,\dots,K$,
$y_{k}\in\{-1,+1\}\ \ \ \mathbb{P}(y_{k}=\pm 1)=\pi_{\pm 1}\ \ \ (X_{k})_{i}|y_{k}\sim N(\mu_{y,k},\Sigma).$ (1)
@@ 2 Transfer Learning via Regularized Discriminant Analysis
Here $(X_{k})_{i},i=1,\cdots,n_{k}$ is a $p-$dimensional vector, and for ease of notation, we write $\mathbf{X}_{k}=((X_{k})_{1}\,(X_{k})_{2}\,\ldots(X_{k})_{n_{k}})^{\top}\in\mathbb{R}^{n_{k}\times p}$ as a $n_{k}\times p$ dimensional matrix. For simplicity of notation, we assume that the mixing proportions $\pi_{\pm 1}$ remain the same across populations. In fact, without much loss of generality, we mainly discuss the simpler case $\pi_{-1}=\pi_{+1}=1/2$. The more general case can be managed in a similar manner without significant technical difficulty (Appendix C). Moreover, in most of the paper, we assume the population covariance matrix are the same for all $K$ populations, and are denoted by $\Sigma$. This assumption is relaxed in Section 6 where we allow each population to have their own covariance matrix $\Sigma_{k}$ for $k=1,\dots,K$.
Under this set up, the Bayes optimal prediction direction for the target population $K$ (Anderson, 1958) is given by,
$d_{Bayes}:=\Sigma^{-1}\delta_{K}\quad\text{where}\quad\bar{\mu}_{K}=\frac{\mu_{+1,K}+\mu_{-1,K}}{2}\quad\text{for}\quad\mu_{\pm 1,K}=\bar{\mu}_{K}\pm\delta_{K},$
@@
and $\delta_{K}=(\mu_{+1,K}-\mu_{-1,K})/2$. The Bayes prediction for a testing data point $x_{0}$ from population $K$ is
@@
where $x_{0}\in\mathbb{R}^{p}$ is a test point in the target population. We call this TL-RDA. We use the regular intercept term $\widehat{b}_{K}$ as all intercept terms are asymptotically zero in the regime considered.
We formulate two criteria to be optimized over $\mathbf{w}$. Let us recall the Bayes classification direction $d_{Bayes}$. Now for a testing data pair $(x_{0},y_{0})$ sampled from population $K$, we consider the two errors:
@@
In addition, we assume that the population covariance matrix are the same for all $K$ populations (assumption 1). With this in mind, a natural way to further exploit mutual information across populations is to use a pooled sample covariance for all discriminant directions $\widehat{d}_{k}$. We write $\widehat{\Sigma}_{P}$ to denote the pooled sample covariance matrix
@@
This form of TL estimator (2) has been used in Zhang and Li (2023) and Dobriban and Sheng (2020) for regression methods. Helm et al. (2024) also considers a TL estimator that aggregates auxiliary information by weighted summations of individual discriminant directions. However, their aggregation is less adaptive in the form of $\widehat{d}(\mathbf{w})^{*}=a\widehat{d}_{1}+(1-a)\sum_{k=2}^{K}\widehat{d}_{k}/(K-1)$ where $a$ is a constant. In addition, the final weight $a$ minimizes only empirical quantities and provides no guarantee of minimizing the population level criteria such as (3) or (4).
3 Random-effects LDA and the Assumptions on Population Parameters
To find the weights that minimize the estimation or prediction risk defined as 3 and 4, we consider random-effects LDA where the population means are random and are potentially correlated across different models. In addition, we also make the random matrix assumptions on the covariance matrix, which allows us to apply the recent advances in random matrix theory to derive the limiting values of the weights and the corresponding classification errors.
3.1 Classification Weights
A key parameter in two-class classification problem is the classification weights $\delta_{k}$ that separate the two classes. In this paper, we consider a random classification weights set up formalized below.
Assumption 3.1 (Random Classification Weights).
The following conditions hold for populations $k=1,\cdots,K$.
  1. 1.
    for fixed $C,\eta>0$.
  2. 2.
    $\bar{\mu}_{k}$ are either fixed or randomly distributed independent of $\delta_{k},\mathbf{X}_{k},y_{k}$. The second moment of $\bar{\mu}_{k}$ are bounded almost surely such that $\lim\sup_{p\rightarrow\infty}\|\bar{\mu}_{k}\|/p^{1/2-\xi}\leq C$ for some constants $\xi,C>0$.
Assumption 3.2 (Correlated Classification Weights).
3.2 Random Matrix Assumption and Related Results
The sample covariance matrix is an important part of the transfer learning estimator. We consider the Marchenko-Pastur type sample covariance matrices as in several relevant works (Dobriban and Wager, 2018; Zhao et al., 2023; Zhang and Li, 2023). For a symmetric matrix, we can characterize its spectral distribution by a cumulative distribution function that places equal point mass on its eigenvalues. Our asymptotic analysis is based on the convergences of spectral distributions of sample covariance matrices, for which the following assumptions are required.
@@ 3.2 Random Matrix Assumption and Related Results
4 Asymptotic Analysis of Weights and Classification Errors
@@ 4.1 Classification Error
Under the two-class Gaussian classification model (Assumption 1), the expected test error of the linear classifier TL-RDA under weight $\mathbf{w}$ in target population $K$ can be written as
@@
Suppose that assumptions 1-3.2 as well as 4.1 and 4.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, we have that the limiting form of $Err(\mathbf{w})$ for a given weight vector $\mathbf{w}\in\mathbb{R}^{K}$ is given by:
@@
The limits of the off-diagonal terms in $\mathcal{A}$ depends on the interplay of sample covariances in separated populations. Similar terms also appear later in the paper when finding the optimal weight vector (see Theorems 4.3, 4.7). Depending on the specific scenario, the term $\mathcal{M}_{kk^{\prime}}$ has explicit expressions, some of which are presented separately in Lemma 5.
@@ 4.2 Minimum Estimation Risk Weight
Suppose that assumptions 1-3.2 as well as 4.1 and 4.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, the weight for minimizing the error in estimating the Bayes optimal discriminator $d_{Bayes}$ is given by:
@@ 4.3 Minimum Prediction Risk Weight
Suppose that Assumptions 1-3.2 as well as 4.1 and 4.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, the weight for minimizing the excess risk, i.e., the error in predicting the class at a random test point $x_{0}$, when compared to the Bayes optimal discriminator $d_{Bayes}$, is given by:
@@ 4.4 Geometric Interpretation
$d_{est}:=\widehat{d}(\mathbf{w}_{E}),\ \ \ d_{err}:=\widehat{d}(\mathbf{w}_{P}).$
Also recall the Bayes discriminant direction $d_{Bayes}$. These discriminant directions are visualized in Figure 1. The blue plane is the space spanned by linear combinations of the local discriminant directions $\{\widehat{d}_{k}\}$, and the top black line stands for the Bayes direction, which is not necessarily in the linear span of $\{\widehat{d}_{k}\}$. Since both the TL-RDA directions $d_{est}$ and $d_{err}$ are linear combinations of $\widehat{d}_{k}$, they both lie on the blue plane and are denoted by colored lines. The direction obtained by minimizing the error in estimating $d_{Bayes}$ is given by $d_{est}$, which is the projection of $d_{Bayes}$ onto the blue plane, denoting ${\rm span}\{\widehat{d}_{k}\}$. The criterion 3 suggests that $d_{est}$ is the minimizer of the OLS loss when fitting $d_{Bayes}$ with linear combinations of $d_{k}$.
@@
Since $\Phi^{-1}(\cdot)$ is a monotonically increasing function, it is clear that $\cos(\theta)$ is close to one, i.e., $\theta$ is close to zero, if and only if $Err(\mathbf{w})$ is close to the Bayes error $Err_{Bayes}$. That is, the size of $\theta$ directly quantifies the inefficiency of TL-RDA relative to $d_{Bayes}$ in terms of classification error: a smaller $\theta$ is equivalent to a near-optimal transfer combination weight $\mathbf{w}$.
Finally, let us also consider an observation $x_{0}$ pointing in a random direction, and denote the angle between $d_{err}$ and $x_{0}$ as $\beta$ and the angle between $d_{Bayes}$ and $x_{0}$ as $\alpha$. Straightforwardly, $d_{err}$ minimizes the difference between $\beta$ and $\alpha$ as $\mathbf{w}^{P}$ minimizes the difference between the inner products $d_{err}^{\top}x_{0}$ and $d_{Bayes}^{\top}x_{0}$. Proposition 4.11 suggests $d_{err}$ minimizes $\theta$ simultaneously.
@@
5 Robustness and Weight Selection
In this section, we present some guidance on how one can choose between different weighting schemes. As we have repeatedly illustrated in the previous section, the $d_{est}$ is dominated by $d_{err}$ in terms of classification error on unseen data. This claim, however, holds only when the test data distribution is as given by Assumption 1. As we shall demonstrate, $d_{est}$ out performs $d_{err}$ considerably when there is a distribution shift in terms of classification error. The TLP-LDA estimator instead consider the space spanned by $\widehat{d}_{k}^{P}$, thus, the final TL estimator is in a different but related column space. We will prove, at least in special cases, optimal TLP-RDA outperforms the optimal TL-RDA when the aspect ratio is large enough.